OpenStudy (anonymous):

The sum of three consecutive numbers is twelve less than six times the lowest number. What are the three numbers?

OpenStudy (aric200):

x+x+1+x+2=6x-12 3x+3=6x-12 15=3x x=5 x+1=6 x+2=7